OpenStudy (anonymous):

ok my cousin gave me this math problem to do and i have no idea what to do!!! anyone know??? 1. One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

i have no idea at all i want to prove him wrong cause he said i cant figure it out so yeah

OpenStudy (arlen):

well horses have got 4 legs and human 2 but both have got 1 head

OpenStudy (anonymous):

true but how to figure it out lol

hartnn (hartnn):

so u just need to set up 2 equations: let x humans and y horses, x+y=78 --->(1) 2x+4y=196---->(2) now solve these,can u ?

OpenStudy (anonymous):

no:(

OpenStudy (arlen):

2x+2y=156 2x+4y=196 2y= 40 y=20

hartnn (hartnn):

ok, from 2nd equation we have , x+2y= 98---->(3) now do (3) - (1)

OpenStudy (anonymous):

2

OpenStudy (arlen):

x+y=78...now y=20 x+20=78 x=78-20=58

hartnn (hartnn):

@arlen solved it for u ,look through it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i see what he did and i looked threw it i asked my cousin if it was right and he said wrong the answer is horses=24 humans=50

OpenStudy (anonymous):

i have no idea how he got that??

hartnn (hartnn):

oh,yes. thats correct.

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

2x+2y=148 2x+4y=196 2y= 48 y=24=horses

hartnn (hartnn):

x+y=74...now y=24 x+24=74 x=74-24=50=humans